OpenStudy (anonymous):

How many grams of the following nonelectrolytes would have to be mixed with 100.0 g of p-dichlorobenzene to increase the boiling point by 3.0 deg C? To decrease the freezing point by 2.0 deg C? a) Succinic Acid b) Caffeine

OpenStudy (aaronq):

Use: \(\huge \sf \Delta T =i*m*\small \underbrace{\huge K_x}_{\small x~=~b~or~f}\) since they're asking for a non-electrolyte, so \(i=1\) \(\large \sf \Delta T =m*\underbrace{K_x}_{\small x~=~b~or~f}\) ------------------------------------------------------------ Make some substitutions with molality and moles, \(\sf m=molality=\dfrac{n_{solute}}{L_{solution}}\) \(\sf n=moles=\dfrac{mass}{Molar~mass}\) ------------------------------------------------------------ \(\large \sf \Delta T =\dfrac{\dfrac{mass}{Molar~mass}}{L_{solution}}* \underbrace{ K_x}_{\small x~=~b~or~f}=\dfrac{mass_{solute}* K_x}{Molar~mass*L_{solution}}\) \(\large \sf mass_{solute}=\dfrac{\Delta T*Molar~mass*L_{solution}}{K_x}\)